OpenStudy (anonymous):

Rationalize: 1/the square root of 2 ?????

OpenStudy (anonymous):



OpenStudy (anonymous):

\[\frac{1}{\sqrt{a}}=\frac{\sqrt{a}}{a}\]

OpenStudy (anonymous):

so it is equal to the square root of 2 over 2?

OpenStudy (ryan123345):